# Chain-of-Discussion: A Multi-Model Framework for Complex Evidence-Based Question Answering

### v1

@@ 6.3 Limitations of Open-Source LLMs
Recalling the results in Table 1 where using multi-model interaction only at one stage fails to enhance Xverse-7B or Qwen-7B. We have to acknowledge that due to the limitations in parameter size, these models might produce unreliable output or demonstrate inferior capabilities to follow instructions.
We guess there can be two reasons. (1) The target LLM may fail to discern errors in the question analysis of other LLMs, and integrate them into the summary. (2) The target LLM does not always follow instructions, and refuse to modify the errors in evidence analysis.
6.3.1 Hallucination Propagation in Summary
At the stage of question analysis, we explicitly instruct the target LLM to assess whether the question analyses of other LLMs are logically correct and relevant to the question. However, the target model might still incorporate hallucinated analyses into the summary.
@@ 6.3.1 Hallucination Propagation in Summary
Table 4: A bad case of hallucinated summary of question analysis. Violet texts are correct analysis, while the texts with yellow background are hallucinated parts.
Table 4 illustrates a bad case where Baichuan-7B incorporates inaccurate analysis from DeepSeek-7B into the summary.We find that LLMs can correctly identify the primary intention behind user’s question, i.e., determining whether providing support for parents is a prerequisite for inheriting. However, Deepseek-7B also engages in analysis regarding scenarios where parents should pay support for children. And Baichuan2-7B fails to identify that this part of analysis is irrelevant to the question.
For people with legal background, it is a commonsense that raising children and caring for elderly parents are two different concepts. And legally, there is no causal relationship between them. However, we guess that for LLMs, both raising children and supporting parents represent financial support and life care between two generations. We believe that open-source LLMs with around 7B parameters may not have sufficient knowledge and enough reasoning capabilities. They may judge whether a question analysis is relevant to the question based on shallow semantic similarities.
@@ 6.3.2 Model Preference of Revising
Similar to the discussion on question analysis, the lack of good reasoning capabilities may also limit the performance of LLMs in revising evidence analysis. We find that LLMs sometimes tend to focus on the literal differences between questions and law articles, illustrated as the original analysis in Table 5. We humans have the background knowledge that only courts have the authority to revoke guardianship, while the LLMs cannot yield correct analysis. Users often omit commonsense information when posing questions, whereas law articles exhibit a rigorous and detailed structure. The stylistic differences between questions and evidence may lead to the misjudgments of LLMs.
Low-quality modification:
High-quality modification:
We also find that different LLMs have varying preferences regarding whether to modify the evidence analysis. For instance, compared to Qwen-7B, Baichuan2-7B exhibits a greater preference to revise the analyses regarding articles.
To quantitatively study this issue, we initially employ Baichuan2-7B to generate the original evidence analyses. Deepseek-7B and Xverse-7B then play the role of critics to determine which analysis should be revised. Finally, we use Baichuan2-7B and Qwen-7B to revise these analyses, respectively. We find that Baichuan2-7B successfully revises 96.5% of the analyses, while Qwen-7B can only revise 56.1% of them.
To mitigate the influence of target model selection on the conclusions, we also use Qwen-7B to provide origianl evidence analyses, still with Deepseek-7B and Xverse-7B as the critics. Similarly, Baichuan2-7B can revise 92.5% of the analyses, but Qwen-7B only revise 67.2% of them.
We argue that an LLM’s preference for refusing to revise may lead to a failure to obtain better evidence analysis based on the critiques. Consequently, it might result in the Chain-of-Discussion framework not bringing enough improvement as expected. The preference of LLMs can be affected by supervised fine-tuning and reward modeling Ouyang et al. (2022); Rafailov et al. (2023). We hope to study the effect of supervised training on Chain-of-Discussion in future.
@@ 7 Conclusions
In this work, we proposed a novel reasoning framework, Chain-of-Discussion, for complex evidence-based question answering tasks. The CoD framework involves multiple LLMs in discussions to achieve more correct and comprehensive responses with less hallucination and more supportive evidence. Experiments on a legal consultation dataset show CoD can effectively improve the performance of open-source LLMs by encouraging them to discuss and criticize.
@@ Appendix A Details of Generation
We note that an input sequence can consist of up to 4K tokens, including the in-context-learning examples. However, we find Qwen-14B Bai et al. (2023) and Baichuan2-13B Baichuan (2023) show suboptimal abilities to process long input sequences. Different from the LLMs with 7B parameters, these models may fail to follow instructions and generate fragment-like analyses. Therefore, we cannot provide the results of Qwen-14B and Baichuan2-13B as baselines.
@@ Appendix C Rules to Identify the Article Reference
Appendix D Prompts of Chain-of-Discussion
D.1 Prompt of Question Analysis
To obtain the question analysis, we employ the prompts as below:
你是一个民法领域的专家,你需要从法律专业的角度分析一名咨询者提出的问题涉及哪些关键点。在分析问题之后,你还要分析检索器提供的参考法条是否能作为分析该问题的依据。请你紧紧围绕咨询者的问题进行分析,不要过度设想潜在的、与问题不相关的场景。
@@ D.1 Prompt of Question Analysis
接下来,请你分析咨询者的问题“[[question]]”
请你站在咨询者的角度,首先分析咨询者所处的环境及其面对的客观事实,再围绕着咨询者的问题,分析他可能需要了解哪些方面的法律规定。你需要先对问题进行分析,再分析各个参考法条是否有该问题有关。请你遵循格式,以“问题分析:”、“法条分析:”分别作为两段分析的开头。
Since we find the LLMs tend to provide the article analysis accompanied with question analysis. We ask it to generate these two analyses separately. We stop the generation after ”法条分析:” has been outputed.
D.2 Prompts of Summary
@@
你是一名法律专家,你需要评价数位律师针对某个法律咨询问题的分析,并给出最终的总结性分析。这些律师的分析既包含正确的内容,也包含错误的内容。你可以参考法条中的内容,谨慎判断各位律师对问题本身的分析是否正确,以及对潜在情况的讨论是否合理。请注意,多数律师都提及的内容更有可能是正确的,你在最终总结时可以参考这部分内容。而如果某个内容仅被个别律师提及,那么该内容有可能是错误的或者与问题不相关的,你需要谨慎判断它是否合理。
@@
D.3 Prompts of Evidence Analysis
@@
你是一名法律专家,你需要判断某个法条是否能作为依据,用于解答给定的法律咨询问题。针对咨询者提出的问题,检索器提供了五个参考法条。但是这些法条可能对于解答问题有帮助,也可能没有帮助。你需要逐个对法条进行分析和判断,在针对某个法条进行判断时,请你不要对其他法条进行判断。
@@
检索器提供的参考法条:
[[articles]]
@@
D.4 Prompt of Critique
When criticizing other LLM’s article analysis, we use following prompts:
你是一名法律专家,你需要点评一名律师对于某个法条是否有助于解答某个法律咨询问题的分析是否误解了法条的内容。我会明确告知你问题和法条的具体内容。
@@ D.4 Prompt of Critique
D.5 Prompt of revising
@@
你是一名律师,你对于某个法条是否有助于解答某个法律咨询问题进行了点评。一些法学专家认为你的点评中存在对法条内容的理解、法条与问题之间的关联性等角度存在错误。你需要参考你对问题的分析,修改你对法条的分析。
@@
D.6 Prompt of response
@@
你是一个法律专家,你需要从法律专业的角度回答咨询者提出的问题。你需要以具体的法条为依据回答问题,并告诉咨询者法律赋予他哪些权利和义务,或者禁止他实施哪些举措。在回答问题之前,你可以参考检索器提供的一些参考法条。但请注意,检索器提供的法条并不一定都有助于回答咨询者提出的问题,它也可能与提问者的问题无关。因此,你需要对问题涉及的事实背景进行分析,再判断各个法条是否能够作为回答问题的依据。请你不要将检索器提供的全部参考法条都当作依据,也不要引用参考法条之外的其他法条作为依据。在回答的过程中,请你紧紧围绕提问者的问题进行讨论,不要过度设想潜在的、与问题不相关的情形。
咨询者的问题是“[[question]]”
下面是检索器提供的参考法条: [[articles]]
接下来,请你回答咨询者提出的问题“[[question]]” 你需要先对该问题的关键点进行分析,然后判断各个参考法条是否有助于解答该问题。最后请你使用与该问题有关的部分法条作为依据,给出详细的回答。回答过程中禁止使用参考法条之外的内容。
问题分析:[[que_ana]]
法条分析:[[art_ana]]
回答:
Appendix E Scoring Prompt of GPT-4

### v2

@@ 6.3 Limitations of Open-Source LLMs
As shown in Table 1, using multi-model interaction at only one stage fails to enhance Xverse-7B (S1) or Qwen-7B (S2). We guess this may be due to the small size of the model scales, 7B here.
After a manual check, we find the main difficulties for those small scaled open-source LLMs: 1) those 7B LLMs may not have sufficient legal/background knowledge and reasoning capabilities, thus often fail to discern errors in the analysis of others and integrate them into the summary; 2) the small scaled LLMs do not always follow instructions, and often refuse to modify the errors in evidence analysis.
@@
Table 3: An example of hallucinated summary of question analysis. Violet texts are correct analysis, while the texts with yellow background are hallucinated parts.
Hallucination Propagation in Summary. Table 3 illustrates an example where Baichuan-7B incorporates inaccurate analysis from DeepSeek-7B into the summary.We find that both LLMs can correctly identify the primary intention from the question, i.e., determining whether providing support for parents is a prerequisite for inheriting. Unfortunately, Deepseek-7B mistakenly engages in analysis regarding whether parents pay support for children, while Baichuan2-7B fails to identify this as irrelevant to the question.
Although both raising children and supporting parents are financial support and life care between two generations in a family, it is a commonsense, for people with legal background, that these are two different concepts and legally, there is no causal relationship between them. We think both LLMs actually rely on shallow semantic similarity to judge if a question analysis is relevant to user’s question.
Capturing Nuance between Related Terms. We guess that open-source LLMs with around 7B parameters may not have sufficient knowledge and enough reasoning capabilities to distinguish related but different terms. Table 4 shows an example where Qwen-7B fails to recognize that {CJK*}UTF8gbsn复婚 (reconcile and remarry) and {CJK*}UTF8gbsn再婚 (remarry with other person) represent different marriage relationships, comply with different articles or provisions, and should not be used interchangeably. However, in our CoD setting, the discussions with other models can, to some extent, have the chance to reduce the hallucination caused by the target model’s poor reasoning capabilities, thus help to reduce logical errors in Qwen-7B’s responses.
Question: What is the difference between resumption of marital relationship ({CJK*}UTF8gbsn复婚) and remarriage with other person ({CJK*}UTF8gbsn再婚)?
Reasoning with Knowledge. Similarly, we find the lack of good reasoning capabilities and knowledge may also limit the performance of LLMs in revising evidence analysis. LLMs tend to focus on the literal differences between questions and law articles. Users often omit commonsense information in their questions, whereas law articles exhibit a rigorous and detailed structure. The stylistic differences between questions and evidence may lead to the misjudgments of LLMs (more examples and analysis in Appendix D.2).
@@ 6.3.1 Model Preference of Revising
We find different LLMs have varying preferences regarding whether to modify the evidence analysis.
To quantitatively study this issue, we initially employ Baichuan2-7B to generate the original evidence analyses. Deepseek-7B and Xverse-7B then play the role of critics to determine which analysis should be revised. Then, we use Baichuan2-7B and Qwen-7B to revise these analyses, respectively. We find that Baichuan2-7B successfully revises 96.5% of the analyses, while Qwen-7B can only revise 56.1% of them.
To mitigate the influence of target model selection, we also use Qwen-7B to provide origianl evidence analyses, still with Deepseek-7B and Xverse-7B as the critics. Similarly, Baichuan2-7B can revise 92.5% of the analyses, but Qwen-7B only revise 67.2% of them.
We argue that an LLM’s preference for refusing to revise may lead to a failure to obtain better evidence analysis based on the critiques. Consequently, it might limit our Chain-of-Discussion framework to bringing more improvement as expected. The preference of LLMs can be affected by supervised fine-tuning and reward modeling Ouyang et al. (2022); Rafailov et al. (2023). We hope to study the effect of supervised training on Chain-of-Discussion in future.
@@ 7 Conclusions
In this work, we proposed a novel reasoning framework, Chain-of-Discussion, for complex evidence-based question answering tasks. The CoD framework involves multiple LLMs in discussions to achieve more correct and comprehensive responses with less hallucination and more supportive evidence. Experiments on a challenging legal consultation dataset show CoD can effectively improve the performance of open-source LLMs by encouraging them to discuss and criticize.
@@
Ethics Statement
In this work, we collect the legal consultation question-answer pairs through Web Search Engine. All these data are publicly accessible on the internet. Additionally, these websites all allow crawlers of search engines to automatically access their web pages.
To protect the privacy of the consultants in real world, we manually check all the instances, and remove those which may indicate the name, address, phone number of a specific users. Note that information such as gender, age, and family composition serves as crucial background in judicial practice. Since this information does not identify a specific individual and does not result in privacy leaks, we do not discard data where the users mention their gender, age, or family composition.
@@ Appendix C Rules to Identify the Article Reference
Appendix D More Results and Analysis
D.1 Lengths of Generated Responses
In Table 5 we provide the average lengths of generated responses from each LLM in different settings. We find that the lengths of responses under different experimental settings can be similar. After checking the correlation coefficient between the difference in response lengths and the difference in GPT-4 scores, we believe that the GPT-4 evaluator does not show preference for longer responses.
Model Avg. Lens. of base Avg. Lens. of CoD
Baichuan2-7B 201.20 200.78
Deepseek-7B 214.29 219.09
Qwen-7B 236.55 259.03
Xverse-7B 236.61 244.60
Table 5: The average lengths of responses generated by different LLMs in baseline setting and CoD setting.
D.2 Examples required Reasoning with Commonsense Knowledge
Low-quality modification:
High-quality modification:
As shown in the original analysis of Table 6, LLMs tend to focus on the literal differences between questions and law articles. We humans have the background knowledge that only courts have the authority to revoke guardianship, thus often omit this information in our questions. But, LLMs often struggle with such differences between the questions and rigorous law articles, thus may not yield correct analysis.
D.3 Significant Test
We apologize that we have not provided significance test. We have to admit that the GPT-4 evaluation is indeed too expensive. A single evaluation of all results under all settings can cost \$150. It is difficult for us to repeat the evaluation multiple times, since it may cost thousands of dollars.
In this work, we use a low temperature when generating. And we set the temperature to 0.0 for GPT-4 evaluation.
We randomly sample 20 examples and employ Baichuan2-7B to repeat the experiments of the baseline setting for 5 times with different seeds. We then use GPT-4 to score the responses of every run. The standard error of the average scores is 0.0245, which is much smaller than the improvement brought by CoD.
We also repeat the experiments on Baichuan2-7B for 5 times under the setting of CoD, using the same subset. The standard error of average scores yielded by GPT-4 is 0.0510, which is also smaller than the improvement brought by CoD.
Appendix E Prompts of Chain-of-Discussion
E.1 Prompt of Question Analysis
To obtain the question analysis, we employ the prompts as below ([]indicates the English translation of the prompts in Chinese):
{CJK*}
UTF8gbsn你是一个民法领域的专家,你需要从法律专业的角度分析一名咨询者提出的问题涉及哪些关键点。在分析问题之后,你还要分析检索器提供的参考法条是否能作为分析该问题的依据。请你紧紧围绕咨询者的问题进行分析,不要过度设想潜在的、与问题不相关的场景。
[ You are an expert in the field of civil law. You need to analyze the key points of a question posed by a consultant from a legal professional’s perspective. After analyzing the question, you should also assess whether the reference articles provided by the retriever can serve as a basis for analyzing the issue. Please focus closely on the consultant’s question and avoid overly imagining potential scenarios unrelated to the issue.]
@@ E.1 Prompt of Question Analysis
[ Below are the reference articles provided by the retriever: [[articles]]]
接下来,请你分析咨询者的问题“[[question]]”请你站在咨询者的角度,首先分析咨询者所处的环境及其面对的客观事实,再围绕着咨询者的问题,分析他可能需要了解哪些方面的法律规定。你需要先对问题进行分析,再分析各个参考法条是否有该问题有关。请你遵循格式,以“问题分析:”、“法条分析:”分别作为两段分析的开头。
[ Next, please analyze the consultant’s question "[[question]]" Please put yourself in the consultant’s shoes, first analyzing the environment they are in and the objective facts they are facing. Then, focusing on the consultant’s question, analyze which aspects of the legal articles they might need to understand. You need to first analyze the question, then assess whether each reference article is relevant to the issue. Please follow the format, starting each section of the analysis with "Question Analysis:" and "Article Analysis:" respectively. ]
Since we find the LLMs tend to provide the article analysis accompanied with question analysis. We ask it to generate these two analyses separately. We stop the generation after "{CJK*}UTF8gbsn法条分析:" ([Article Analysis]) has been output.
E.2 Prompts of Summary
@@
{CJK*}
UTF8gbsn你是一名法律专家,你需要评价数位律师针对某个法律咨询问题的分析,并给出最终的总结性分析。这些律师的分析既包含正确的内容,也包含错误的内容。你可以参考法条中的内容,谨慎判断各位律师对问题本身的分析是否正确,以及对潜在情况的讨论是否合理。请注意,多数律师都提及的内容更有可能是正确的,你在最终总结时可以参考这部分内容。而如果某个内容仅被个别律师提及,那么该内容有可能是错误的或者与问题不相关的,你需要谨慎判断它是否合理。
[ You are a legal expert. You need to evaluate several lawyers’ analyses of a specific legal consultation question and provide a final summary analysis. These lawyers’ analyses contain both correct and incorrect elements. You can refer to the content of the legal articles, carefully judging whether each lawyer’s analysis of the issue itself is correct and whether their discussion of potential scenarios is reasonable. Note that content mentioned by most lawyers is more likely to be correct, and you can reference this part in your final summary. If a particular point is mentioned by only a few lawyers, it may be incorrect or irrelevant to the issue, so you need to carefully judge its reasonableness. ]
@@
[ You first need to critique the lawyers’ analyses, evaluating whether each point in these analyses is correct and relevant to the issue. For points with logical errors or low relevance to the issue, you should clearly point them out and provide criticism. Afterward, based on your critiques, provide a summary analysis that matches the tone and style of the lawyers’ analyses of the issue. ]
E.3 Prompts of Evidence Analysis
@@
{CJK*}
UTF8gbsn 你是一名法律专家,你需要判断某个法条是否能作为依据,用于解答给定的法律咨询问题。针对咨询者提出的问题,检索器提供了五个参考法条。但是这些法条可能对于解答问题有帮助,也可能没有帮助。你需要逐个对法条进行分析和判断,在针对某个法条进行判断时,请你不要对其他法条进行判断。
[ You are a legal expert tasked with determining whether a specific legal article can serve as a basis for answering a given legal consultation question. The retriever has provided five reference articles for the question posed by the consultant. However, these articles may or may not be helpful in answering the question. You need to analyze and evaluate each article individually, refraining from making judgments about other articles when assessing a particular one. ]
@@
[ The consultant’s question is:[[question]]]
检索器提供的参考法条: [[articles]]
[ The reference articles provided by the retriever are: [[articles]]]
@@
[ You need to first analyze the key points of the question and then proceed to analyze each article to determine its relevance to answering the question. During the analysis of the articles, consider what rights and obligations the article specifies, or what behaviors it prohibits. If the articles are somewhat relevant to the key points of the question, then the article may likely be helpful in answering the question; otherwise, the article is probably not helpful in addressing the question.]
E.4 Prompt of Critique
When criticize other LLM’s article analysis, we use following prompts:
{CJK*}
UTF8gbsn你是一名法律专家,你需要点评一名律师对于某个法条是否有助于解答某个法律咨询问题的分析是否误解了法条的内容。我会明确告知你问题和法条的具体内容。
[ You are a legal expert. You need to critique a lawyer’s analysis of whether a specific legal article is helpful in answering a particular legal consultation question and determine if their analysis misconstrues the content of the legal article. I will provide you with the specific details of the question and the legal article. ]
@@ E.4 Prompt of Critique
[ Next, please provide a concise critique of the lawyer’s analysis of [[cur_art_id]]. Then, determine whether their analysis misconstrues the content of the legal article. ]
E.5 Prompt of revising
@@
{CJK*}
UTF8gbsn你是一名律师,你对于某个法条是否有助于解答某个法律咨询问题进行了点评。一些法学专家认为你的点评中存在对法条内容的理解、法条与问题之间的关联性等角度存在错误。你需要参考你对问题的分析,修改你对法条的分析。
[ You are a lawyer who has provided an assessment of whether a certain legal article is helpful in addressing a specific legal consultation question. Some legal experts believe there are errors in your assessment regarding understanding the content of the legal article and its relevance to the question. You need to revise your analysis of the legal article based on your analysis of the issue. ]
@@
[ Next, please rewrite a more accurate analysis of the legal article. At the end of the rewritten analysis of the legal article, evaluate whether the legal article may indeed be helpful in addressing the question based on your analysis. ]
E.6 Prompt of response
@@
{CJK*}
UTF8gbsn 你是一个法律专家,你需要从法律专业的角度回答咨询者提出的问题。你需要以具体的法条为依据回答问题,并告诉咨询者法律赋予他哪些权利和义务,或者禁止他实施哪些举措。在回答问题之前,你可以参考检索器提供的一些参考法条。但请注意,检索器提供的法条并不一定都有助于回答咨询者提出的问题,它也可能与提问者的问题无关。因此,你需要对问题涉及的事实背景进行分析,再判断各个法条是否能够作为回答问题的依据。请你不要将检索器提供的全部参考法条都当作依据,也不要引用参考法条之外的其他法条作为依据。在回答的过程中,请你紧紧围绕提问者的问题进行讨论,不要过度设想潜在的、与问题不相关的情形。
{CJK*}
UTF8gbsn[ You are a legal expert, and you need to answer the question posed by the consultant from a legal perspective. You are required to provide specific legal articles as the basis for your answer, informing the consultant of their rights, obligations conferred by the law, or actions prohibited by it. Before answering the question, you can refer to some reference articles provided by the retriever. However, please note that the articles provided by the retriever may not necessarily be helpful in answering the consultant’s question; they may also be irrelevant to the question. Therefore, you need to analyze the factual background of the issue involved, then determine whether each article can serve as a basis for answering the question. Please do not consider all the reference articles provided by the retriever as the basis, nor cite any articles outside the reference ones as evidence. During your response, focus closely on the consultant’s question, avoiding overly imagining potential scenarios unrelated to the issue.]
{CJK*}
UTF8gbsn 咨询者的问题是“[[question]]”
{CJK*}
UTF8gbsn[ The consultant’s question is: "[[question]]"]
{CJK*}
UTF8gbsn 下面是检索器提供的参考法条: [[articles]]
{CJK*}
UTF8gbsn[ Below are the reference articles provided by the retriever: [[articles]]]
{CJK*}
UTF8gbsn 接下来,请你回答咨询者提出的问题“[[question]]” 你需要先对该问题的关键点进行分析,然后判断各个参考法条是否有助于解答该问题。最后请你使用与该问题有关的部分法条作为依据,给出详细的回答。回答过程中禁止使用参考法条之外的内容。
{CJK*}
UTF8gbsn[ Next, please answer the question posed by the consultant "[[question]]" You need to analyze the key points of the question first, then determine whether each reference article is helpful in answering it. Finally, please provide a detailed answer using relevant portions of the articles as the basis. Use of content outside the reference articles is prohibited during the response. ]
{CJK*}
UTF8gbsn 问题分析:[[que_ana]]
{CJK*}
UTF8gbsn[ Question Analysis: [[que_ana]]]
{CJK*}
UTF8gbsn 法条分析:[[art_ana]]
{CJK*}
UTF8gbsn[ Article Analysis: [[art_ana]]]
{CJK*}
UTF8gbsn 回答:
{CJK*}
UTF8gbsn[ Answer:]
Appendix F Scoring Prompt of GPT-4